OpenStudy (anonymous):

Find the slope of the line passing through the points A( 4, 8) and B(6, -3)

OpenStudy (saifoo.khan):

-3 -8 ------ 6 - 4

OpenStudy (saifoo.khan):

-11/2

OpenStudy (anonymous):

\[slope = \frac{y_{2}-y_{1}}{x_{2}-x_{1}} = \frac{-3-8}{6-4} = \frac{-11}{2}\]